### v2

# On the existence of unbiased resilient estimators in discrete quantum systems

Abstract
Cramér-Rao constitutes a crucial lower bound for the mean squared error of an estimator in frequentist parameter estimation, albeit paradoxically demanding highly accurate prior knowledge of the parameter to be estimated. Indeed, this information is needed to construct the optimal unbiased estimator, which is highly dependent on the parameter. Conversely, Bhattacharyya bounds result in a more resilient estimation about prior accuracy by imposing additional constraints on the estimator. Initially, we conduct a quantitative comparison of the performance between Cramér-Rao and Bhattacharyya bounds when faced with less-than-ideal prior knowledge of the parameter. Furthermore, we demonstrate that the $n^{th}$order classical and quantum Bhattacharyya bounds cannot be computed –given the absence of estimators satisfying the constraints– under specific conditions tied to the dimension $m$ of the discrete system. Intriguingly, for a system with the same dimension $m$, the maximum non-trivial order $n$ is $m-1$ in the classical case, while in the quantum realm, it extends to $m(m+1)/2-1$. Consequently, for a given system dimension, one can construct estimators in quantum systems that exhibit increased robustness to prior ignorance.
I Introduction
Quantum metrology studies the limits in precision by estimating physical parameters encoded in quantum systems. Quantum properties such as entanglement and squeezing are used to surpass classical estimation strategies [1, 2, 3]. Applications like quantum phase estimation [4, 5, 6, 7], quantum-enhanced position and velocity estimation [8, 9, 10], quantum illumination [11, 12, 13], quantum thermometry [14, 15, 16, 17], and channel discrimination [18] among others have boosted the field. Parameter estimation gives the theoretical framework to establish the optimal performance of particular estimation protocols and its main figure of merit is the Cramér-Rao bound (CRB) [19, 20].
Paradoxically, to achieve the ultimate precision established by the CRB an infinite number of measurements have to be performed or the prior value of the parameter has to be known. Both scenarios are unlikely or experimentally unfeasible. In the finite sample regime, CRB is frequently unattainable [21, 22, 23]. Several bounds as Hammersley–Chapman–Robbins bound [24], the family of Bhattacharyya bounds (BhBB) [25] or Barankin [26, 27] bounds, and more generally Abel bounds [28], which include the previous bounds as special cases, were introduced. These bounds succeed in establishing a sensible limit for the achievable precision [29], especially in noisy estimation problems [27].
Recently, there has been interest in developing tighter bounds for quantum parameter estimation tasks. Gessner et al. [30] derived lower bounds on the variance of estimators in quantum metrology by increasing the unbiasedness constraints of the estimator, i.e., in the same spirit as their classical analogs were obtained. They presented a hierarchy of increasingly tight bounds that include the quantum CRB (QCRB) [31] at the lowest order and quantum Barankin, Bhattacharyya, and Abel bounds as other particular cases.
In this article, we study the classical and the quantum Bhattacharyya bounds (QBhBB). We compare the performance of these bounds with the CRB when the prior knowledge of the parameter is not sharp deriving a lower bound for the mean squared error (MSE). We illustrate this result for the Mach-Zehnder interferometer by comparing the actual MSE of the estimator with BhB and CRB and we see a clear advantage when using the BhB. Finally, we address the impossibility of computing classical BhB under certain conditions when dealing with discrete probability distributions. We prove the necessary and sufficient conditions for this bound to exist and we show how these conditions get relaxed for the QBhB.
This article is organized as follows: in sec. II, we introduce estimation theory, present the BhB and motivate them using the Mach-Zehnder interferometer setting. In sec. III, we study the cases when the BhB cannot be computed, i.e. the bound is divergent for both, classical probability distributions and quantum systems. Finally, in sec. IV, the conclusions are presented.
II Estimation theory
To extract information from a physical system measurements have to be performed. The measurement outcome is modeled as a random variable following a probability distribution that depends on the state of the system. The information is encoded in the random variable and estimation theory is the branch of statistics that deals with the optimal protocols and the ultimate limit for the information extraction. The CRB establishes a limit on the precision while estimating a magnitude or a parameter. Consider that the outcome of the experiment, denoted by the random variable $X$, follows the family of probability distribution $P_{\theta}(x)$, that is parameterized by the parameter $\theta$ which is being estimated. An estimator $\tilde{\Theta}(x)$ is a function of the random measurement outcome that estimates the value of the parameter $\theta$. The performance of an estimator is characterized by its bias $b(\theta)$, mean squared error (MSE) and variance $(\Delta\tilde{\Theta})^{2}$,
[…]
$(\Delta\tilde{\Theta}_{\theta})^{2}\geq\frac{1}{F_{C}},$ (4)
where $F_{C}$ is the classical Fisher information and is given by
[…]
where $X_{+}$ is the space of events $x$ with nonzero probability $P_{\theta}(x)>0$.
The CRB is always saturable in the asymptotic limit or when the true value of $\theta$ is known [32]. This is not the case in realistic scenarios. In quantum parameter estimation [1], a family of quantum states $\rho_{\theta}$ parametrized by a parameter $\theta$ plus a particular POVM measurement $\{\Pi_{x}\}$ substitute the family of probability distributions, $P_{\theta}(x)$. The lower bound on the MSE of the estimator is now determined by the quantum Cramér-Rao bound (QCRB) [1, 31, 14], its version for locally unbiased estimators reads
$(\Delta\tilde{\Theta}_{\theta})^{2}\geq\frac{1}{F_{Q}},$ (6)
where $F_{Q}=\operatorname{Tr}[\rho_{\theta}L^{2}]$ is the quantum Fisher information. The symmetric logarithmic derivative (SLD) is the operator $L$ satisfying
$\frac{d\rho_{\theta}}{d\theta}=\frac{\rho_{\theta}L+L\rho_{\theta}}{2}.$ (7)
From Eq.(7) it is deduced that the SLD is a hermitian operator and its eigenvectors can form an orthonormal basis. In the derivation of the QCRB, a general POVM $\{\Pi_{x}\}$ was considered but it was shown that a projection on the basis of eigenvectors of $L$ is the optimal measurement that saturates the QCRB [14]. As in the classical estimation theory, the QCRB is always saturable in the asymptotic regime. Furthermore, it has recently been shown to be non-saturable within a Bayesian perspective [33, 23].
II.1 Bhattacharyya bounds and their motivation
Bhattacharyya bounds were introduced in the context of classical parameter estimation problems to give a lower bound for the variance of the estimate [25]. The BhB gives a tighter bound than the CRB by including higher-order derivatives of the probability distribution in the calculation. It is shown that BhB converge to the variance of the best unbiased estimator –the estimator that is unbiased for any point in the region of the allowed values of the parameter–, when the sampling distribution is a member of an exponential family of distributions [34]. The $n^{th}$order BhB bound gives the lowest value of the variance for an estimator that fulfills $\frac{d^{i}b(\theta)}{d\theta^{i}}|_{\theta=\theta_{0}}=0$ for $1\leq i\leq n$. This estimator is unbiased in a larger region than the one obtained for the CRB, as will be shown later, and its variance is closer to the real variance of a reasonable parameter estimation problem, namely where the real value of the parameter is unknown, and is given by [30],
$(\Delta\tilde{\Theta}_{\theta_{0}})^{2}\geq\max_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}\bm{C}\bm{a}}=\bm{\lambda}^{\top}\bm{C}^{-1}\bm{\lambda}.$ (8)
The last equality is valid only when $\bm{C}^{-1}$ exists; whereas if this is not fulfilled, one needs to resort to the maximization problem. The vectors $\bm{\lambda}$ and $\bm{a}$ have $n$ entries and $\bm{C}$ is a $n\times n$ matrix, where $n$ is the order of the bound we are computing. They are defined as [30]
$\bm{\lambda}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix},\>\>\>\>\>\>\>\bm{C}_{kl}=\int_{x\in X_{+}}\frac{\partial^{k}P_{\theta_{0}}(x)\partial^{l}P_{\theta_{0}}(x)}{P_{\theta_{0}}(x)},$ (13)
where
$\partial^{k}P_{\theta_{0}}(x)=\frac{\partial^{k}P_{\theta}(x)}{\partial\theta^{k}}\bigg{|}_{\theta=\theta_{0}}.$
We would like to point out the similarities between the $n^{th}$order BhB bound and the multiparameter Cramér-Rao bound (MCRB) [35] in a multiparameter estimation problem with $n$ parameters. The computational effort to compute both bounds is the same and the BhB can be codified as a particular case of the CRB in a multiparameter estimation setting. The BhB can be extended to quantum estimation theory [30] in a coherent way that includes them as a specific case of a larger set of bounds. Doing so, the QBhB state that the variance of any unbiased estimator with null $n$ first bias derivatives satisfies
$(\Delta\hat{\Theta}_{\theta_{0}})^{2}\geq\max_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}\bm{Q}\bm{a}}=\bm{\lambda}^{\top}\bm{Q}^{-1}\bm{\lambda}.$ (14)
Again, the last equality holds when $\bm{Q}^{-1}$ exists. The $n\times n$ matrix $\bm{Q}$ is defined as
$\bm{Q}_{kl}=\operatorname{Tr}\left(\frac{d^{k}\rho_{k}}{d\theta^{k}}L_{l}\right),$ (15)
where $L_{l}$ is the generalization of the symmetric logarithmic derivative that satisfies
$\frac{d^{l}\rho_{\theta}}{d\theta^{l}}=\frac{\rho_{\theta}L_{l}+L_{l}\rho_{\theta}}{2}.$ (16)
II.2 An example: Bhattacharyya bound in Mach-Zehnder interferometer
Historically more attention has been paid to the Barakin bounds [26, 27]. The Bhattacharyya bounds have not been computed in any physical problem before. Here, we compute BhB and the estimators achieving the bound for the paradigmatic Mach-Zehnder interferometer. We show that when the prior is not sharp, and we move away from the point for which the unbiased estimators were constructed, the BhB bound offers an improvement to the achievable precision.
$\varphi_{a}$$\varphi_{b}$$a$$b$$a^{\prime}$$b^{\prime}$|$\ket{r}$|$\ket{r}$$n_{a}$$n_{b}$
Figure 1: Mach-Zehnder interferometer setup
II.2.1 Mach-Zehnder
In the Mach-Zehnder configuration, depicted in Fig. (1), two modes of the electromagnetic field interfere on a balanced beamsplitter, the output beams acquire a relative phase $\theta=\varphi_{a}-\varphi_{b}$, and interfere again on the second beamsplitter. Finally, the photon numbers $n_{a}$ and $n_{b}$ are measured at the output ports. We consider the input state |$\ket{r,r}$ where both input modes are in the Fock state with $r$ photons. It has been shown that this state achieves the Heisenberg limit [36] while estimating the relative phase $\theta$. The evolution induced by the interferometer is given by the unitary $U=e^{\frac{\theta}{2}(a^{\dagger}b-b^{\dagger}a)}$ [6]. The probability distribution $P_{\theta}(2q)$ is then
$P_{\theta}(2q)=|\langle r-q,r+q|e^{\frac{\theta}{2}(a^{\dagger}b-b^{\dagger}a)}|r,r\rangle|^{2}$ (17)
where $2q=n_{a}-n_{b}$ is the photon counting difference. The width of this probability distribution is $|q|<r\theta$. If $\theta\ll 1$ the probability distribution can be well approximated by [36]
$P_{\theta}(2q)=J^{2}_{q}(r\theta),$ (18)
[…]
II.3 Analytical motivation
In this section, we derive a lower bound for the MSE and show that it depends on the bias. We first define $\hat{\Theta}_{\theta_{0}}(x)$ as an estimator unbiased at $\theta_{0}$. We compute the MSE of this estimator in the point $\theta^{\prime}$
[…]
We now use the Cauchy-Schwarz inequality to find a lower bound for Eq. (19). Doing so, one gets
[…]
Making the choice $f(x)=(\theta^{\prime}-\tilde{\Theta}_{\theta_{0}})\sqrt{P_{\theta^{\prime}}(x)}$ and $g(x)=\sqrt{P_{\theta^{\prime}}(x)}$, we obtain
[…]
The first integral is the square of the bias, the second is the MSE, and the last integral is just the unity. So we can conclude that
$b^{2}(\theta^{\prime})\leq\text{MSE}(\theta^{\prime}).$ (22)
The squared bias gives a lower bound for the MSE. Since the derivatives of the bias for the BhB estimator are zero, this estimator will have a smaller lower bound for the MSE for an interval $\theta_{0}-\epsilon<\theta^{\prime}<\theta_{0}+\epsilon$. Unfortunately, the saturability of (22) highly depends on the particular probability distribution and it is not possible in general. Nevertheless, in Appendix (A) we prove that if the probability distribution satisfies certain conditions there always exists an interval where the $n^{th}$order BhB performs at least as well as the CRB.
III Conditions of the existence of estimators
Up to this point, we presented the usefulness of the BhB for experimentally friendly scenarios. However, one cannot always construct estimators that fulfill the conditions needed to lower bound their MSE by the BhB. This is true especially if one deals with discrete probability distributions. In this section, we give the conditions for the existence of the estimators for finite support probability distributions. We conclude that in some cases, higher-order BhB might not be computable or irrelevant. The largest BhB computable – and that gives us new information– is related to the dimension of the support of the probability distribution. We also show that the BhB diverges if and only if no estimator is satisfying the unbiasedness conditions.
We present our results for both classical and quantum parameter estimation settings. The problem of quantum parameter estimation encompasses two key optimization challenges: (i) the optimization across all conceivable observables or, more broadly, the optimization across all feasible POVMs that can be executed; and (ii) the optimization across all potential estimators $\tilde{\Theta}$ that can be derived from the outcomes of measurements. While optimization (ii) is addressed by the classical Cramér-Rao bound, optimization (i) introduces a distinctly quantum dimension to the problem. We will demonstrate the implications of this extra freedom to select a POVM.
III.1 Classical
We consider a family of N-point support probability functions $P_{\theta}(x)$ where $P_{\theta}(x_{i})\geq 0$ for $x_{i}\in\{x_{1},..,x_{N}\}$. The family of probability distributions is parametrized by the parameter $\theta$ that we would like to infer.
The most general way to write an estimator $\tilde{\Theta}(x)$ is,
$\tilde{\Theta}(x)=\begin{cases}\tilde{\theta}_{1}&\quad\text{if }x=x_{1},\\ \tilde{\theta}_{2}&\quad\text{if }x=x_{2},\\ \vdots&\quad\vdots\\ \tilde{\theta}_{N}&\quad\text{if }x=x_{N}.\\ \end{cases}.$
We want to compute an unbiased estimator in the point $\theta_{0}$ and impose the first $n$ derivatives of the bias to be zero in the point $\theta_{0}$. Given that $b(\theta)=\langle\tilde{\Theta}\rangle_{\theta}-\theta$ the conditions are
$\begin{split}\langle\tilde{\Theta}\rangle_{\theta_{0}}&=\theta_{0},\\ \frac{d\langle\tilde{\Theta}\rangle_{\theta_{0}}}{d\theta}\bigg{|}_{\theta=\theta_{0}}&=1,\\ \frac{d^{l}\langle\tilde{\Theta}\rangle_{\theta_{0}}}{d\theta^{l}}\bigg{|}_{\theta=\theta_{0}}&=0.\>\>\>\>\>\>\>\>\>l=2,3,...,n.\\ \end{split}$ (23)
Since we are considering discrete probability distributions the mean value can be written as finite sum $\langle\tilde{\Theta}\rangle_{\theta_{0}}=\sum_{i}P_{\theta_{0}}(x_{i})\tilde{\Theta}(x_{i})$. Defining $\tilde{\Theta}(x_{i})=\tilde{\theta_{i}}$, Eq. (23) can be written as a system of equations $A\vec{x}=\vec{b}$ where
$A=\begin{pmatrix}P_{\theta_{0}}(x_{1})&P_{\theta_{0}}(x_{2})&\dots&1-\sum_{i}^{N-1}P_{\theta_{0}}(x_{i})\\ \partial P_{\theta_{0}}(x_{1})&\partial P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})\\ \partial^{2}P_{\theta_{0}}(x_{1})&\partial^{2}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \partial^{n}P_{\theta_{0}}(x_{1})&\partial^{n}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{n}P_{\theta_{0}}(x_{i})\\ \\ \end{pmatrix},\quad\quad\vec{x}=\begin{pmatrix}\tilde{\theta}_{1}\\ \tilde{\theta}_{2}\\ \vdots\\ \tilde{\theta}_{N}\end{pmatrix},\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix},$ (24)
where we defined $\partial^{l}P_{\theta_{0}}(x)\equiv\frac{\partial^{l}P_{\theta}(x)}{\partial\theta^{l}}|_{\theta=\theta_{0}}$. The system has $n+1$ equations and $N$ variables. We consider the case where the first $N-1$ derivatives are linearly independent. This implies that if $n\leq N-1$, there is always a solution. If $n>N-1$, then the higher derivatives are no longer independent, and they can be written as a combination of the first $N-1$ linearly independent derivatives. In this case, the system has no solution or the solution is trivial, that is to say, the $n^{th}$order BhB is equal to $N-1$ order BhB for all $n>N-1$. It is then concluded that given a $N$ points probability distribution is advisable to compute BhB bound up to order $N-1$.
In the appendix (B), the conditions for the existence of the estimator are given. It is also shown that the bound diverges if and only if the estimator does not exist.
III.2 Quantum
We extend the previous results to determine when computing QBhB gives more information about the estimation problem. In quantum estimation theory the measurement is a POVM $\{\Pi_{i}\}$, where $i$ are the possible outcomes, each one with probability $P_{\theta}(i)=\operatorname{Tr}(\Pi_{i}\rho_{\theta})$. The classical processing of the measurement outcomes is accounted by the estimator $\tilde{\Theta}(x)$ that associates an estimate $\tilde{\Theta}(i)$ to the outcome $i$. Every estimate $\tilde{\Theta}(i)$ has a probability given by $P_{\theta}(i)$. Both tasks, measurement, and classical processing, can be treated simultaneously as a hermitian operator $\hat{\Theta}=\sum_{i}\tilde{\Theta}(i)\Pi_{i}$ [22]. The conditions of Eq. (23) are now written in terms of the operator $\hat{\Theta}$ as
$\begin{pmatrix}\operatorname{Tr}[\rho_{\theta_{0}}\hat{\Theta}]\\ \operatorname{Tr}[d\rho_{\theta_{0}}\hat{\Theta}]\\ \operatorname{Tr}[d^{2}\rho_{\theta_{0}}\hat{\Theta}]\\ \vdots\\ \operatorname{Tr}[d^{n}\rho_{\theta_{0}}\hat{\Theta}]\\ \end{pmatrix}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\\ \end{pmatrix},$ (35)
where again
$d^{k}\rho_{\theta_{0}}=\frac{d^{k}\rho_{\theta}}{d\theta^{k}}\bigg{|}_{\theta=\theta_{0}}.$
We rewrite these conditions as
$\begin{pmatrix}(\rho_{\theta_{0}})_{ij}\hat{\Theta}_{ji}\\ (d\rho_{\theta_{0}})_{ij}\hat{\Theta}_{ji}\\ (d^{2}\rho_{\theta_{0}})_{ij}\hat{\Theta}_{ji}\\ \vdots\\ (d^{n}\rho_{\theta_{0}})_{ij}\hat{\Theta}_{ji}\\ \end{pmatrix}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\\ \end{pmatrix},$ (46)
where we are using the Einstein summation convention and $i,j=1,2,...,N$. Now we have a system of equations with complex coefficients for the complex variables $\hat{\Theta}_{ij}$. Since the estimator operator must be hermitian, not all the variables $\hat{\Theta}_{ij}$ are independent. Instead of $N^{2}$ independent variables, there are now $\frac{N(N+1)}{2}$ independent variables. In that way, we can vectorize the system of equations as follows
$\begin{pmatrix}(\rho_{\theta_{0}})_{1}&(\rho_{\theta_{0}})_{2}&\dots&(\rho_{\theta_{0}})_{\frac{N(N+1)}{2}}\\ (d\rho_{\theta_{0}})_{1}&(d\rho_{\theta_{0}})_{2}&\dots&(d\rho_{\theta_{0}})_{\frac{N(N+1)}{2}}\\ (d^{2}\rho_{\theta})_{1}&(d^{2}\rho_{\theta})_{2}&\dots&(d^{2}\rho_{\theta})_{\frac{N(N+1)}{2}}\\ \vdots&\vdots&\ddots&\vdots\\ (d^{n}\rho_{\theta})_{1}&(d^{n}\rho_{\theta})_{2}&\dots&(d^{n}\rho_{\theta})_{\frac{N(N+1)}{2}}\\ \end{pmatrix}\begin{pmatrix}\hat{\Theta}_{1}\\ \hat{\Theta}_{2}\\ \vdots\\ \hat{\Theta}_{\frac{N(N+1))}{2}}\end{pmatrix}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix}.$ (47)
So we have a system of equations equivalent to the one for a classical probability distribution with $\frac{N(N+1)}{2}$ support points. As before if this system of equations cannot be solved the QBhB variance will go to infinity. On the other hand, if the first $\frac{N(N+1)}{2}-1$ derivatives are linearly independent, for all $n>\frac{N(N+1)}{2}-1$ one has that $n^{th}$order QBhB is equal to $\frac{N(N+1)}{2}-1$ order QBhB. The difference between the quantum and classical cases is that in the former we have the freedom to choose a measurement. This extra freedom implies that the highest relevant QBhB bound –which gives a different bound than lower orders– is lifted in the quantum case because $N-1<\frac{N(N+1)}{2}-1$.
Let us elaborate more on this because at first glance the result may seem to be at odds with the classical case. After projective measurements, one deals with $N$ point probability distributions. Therefore, the results derived in the classical setting must also apply, so it might seem that there will be a contradiction to what we stated here. To solve this apparent contradiction, we explain why $N$ order QBh bound need not be equal to the $N-1$ one. The key point is that to compute the $N$ and the $N+1$ order QBhB, different measurements have to be performed, giving different probability distributions $P_{N}$ and $P_{N+1}$. This extra freedom in the measurements is what gives the quantum systems more non-trivial information, or in other words, the maximal order for which QBhB gives different information is higher than for classical scenarios. Thus, since $P_{N}$ and $P_{N+1}$ are different probability distributions, there is no discrepancy with what we presented in section (III.1).
Figure 2: QCR and QBhB as a function of $\lambda$. One sees that they differ and as expected QBhB bound will be higher than QCR. In addition, we plot these values of $\lambda$ because the plot is symmetric around $\lambda=0.5$ and also divergent for this same value –since the state will be $\rho\propto\mathbb{I}$ and won’t contain any information on the parameter–.
III.3 Example of the quantum case
In this section, we show a case in which we see the lifting in the constraints produced by the quantum nature of the problem, meaning that we see that the quantum bounds are not trivial –giving the same value as bounds from lower orders– for $n>N-1$, being $n$ the order of the BhB employed and $N$ the number of outcomes.
Consider that our state $\rho(\theta)$ is obtained through the following unitary operation $U=e^{-i\theta^{2}H}$ where $H=\sigma_{x}$ is the $x$-Pauli matrix, and we would like to estimate $\theta$. The number of outcomes here is $N=2$. Our initial state is given by
$\rho(0)=\frac{1}{2}\mathbb{I}+\frac{(2\lambda-1)}{2}\sigma_{z},$ (48)
where $\lambda\in(0,1)$ and
$\rho(\theta)=U\rho(0)U^{\dagger}=\left(\begin{array}[]{cc}\left(\lambda-\frac{1}{2}\right)\cos\left(2\theta^{2}\right)+\frac{1}{2}&\frac{1}{2}i(2\lambda-1)\sin\left(2\theta^{2}\right)\\ i(1-2\lambda)\sin\left(\theta^{2}\right)\cos\left(\theta^{2}\right)&\frac{1}{2}\left((1-2\lambda)\cos\left(2\theta^{2}\right)+1\right)\\ \end{array}\right).$ (49)
We can proceed and calculate $L_{1}$, and $L_{2}$ following the definitions presented in previous sections. Doing some algebra one ends up with the QBhB matrix which reads
$\bm{Q}_{BhB^{2}}=\left(\begin{array}[]{cc}16\theta^{2}(1-2\lambda)^{2}&16\theta(1-2\lambda)^{2}\\ 16\theta(1-2\lambda)^{2}&\frac{16(1-2\lambda)^{2}\left((\lambda-1)\lambda-4\theta^{4}\right)}{(\lambda-1)\lambda}\\ \end{array}\right),$ (50)
where the first element corresponds to the QFI. To finally get the BhB, we take the inverse of the matrix and look at the first element of it. In Fig. (2) we can see the lower bound on the variance of the estimator when using the QCR bound and QBhB. Thus, computing the quantum Bhattacharyya bound gives new non-trivial information even though $n>N-1$.
IV Conclusions
Since long ago, it has been clear in estimation theory that while the CRB is extremely useful and illustrative is not always meaningful. For that reason, different bounds were constructed and introduced in the classical regime. On the other hand, until very recently, there were no works extending these bounds to the quantum regime [30].
In this article, we first motivated the use of Bhattacharyya bounds using a simple example. Essentially, when one does not know the prior value of the parameter exactly, but knows that is within a small range, the BhB are more desirable. We then gave necessary and sufficient conditions for an estimator that satisfies the BhB to exist. The existence conditions are related to the support of the probability distribution which determines the freedom we have to construct an estimator. When we computed the $n$-th order BhB bound we found a limit for the variance of an estimator that satisfies $n+1$ unbiasedness conditions. When the support of the probability distribution is small, we might not have enough freedom to construct an estimator satisfying the desired constraints. Namely, the estimator exists if and only if the maximization presented in Eq. (8) can be calculated (non-divergent). Moreover, following a similar reasoning, we could anticipate when computing higher orders in the BhB gives no extra information. Specifically, we saw that for $N$ outcomes, it is “sufficient” to calculate the $n^{th}$order BhB bound where $n\leq N-1$. However, in the quantum case, where we have more freedom in the measurements, the order is lifted to $n\leq N(N+1)/2-1$ implying that in the quantum regime, the estimators are more resilient. Finally, it would be of theoretical and practical interest to extend the results presented here to continuous probability distributions as well as to biased estimators. For the latter case, we think it will be conceptually difficult to compare the different bounds such as the BhB and CRB.
Acknowledgements.
We acknowledge fruitful discussions with Manuel Gessner and Maximilian Reichert. The authors acknowledge financial support from OpenSuperQ+100 (Grant No. 101113946) of the EU Flagship on Quantum Technologies, as well as from the EU FET-Open project EPIQUS (Grant No. 899368), also from Project Grant No. PID2021-125823NA-I00 595 and Spanish Ramón y Cajal Grant No. RYC-2020-030503-I funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future,” this project has also received support from the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call –Quantum Spain, and by the EU through the Recovery, Transformation and Resilience Plan– NextGenerationEU within the framework of the Digital Spain 2026 Agenda, we acknowledge funding from Basque Government through Grant No. IT1470-22 and the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government.
Appendix B Existence of the estimator
We study the solution of the system $A\vec{x}=\vec{b}$ with
$A=\begin{pmatrix}P_{\theta_{0}}(x_{1})&P_{\theta_{0}}(x_{2})&\dots&1-\sum_{i}^{N-1}P_{\theta_{0}}(x_{i})\\ \partial P_{\theta_{0}}(x_{1})&\partial P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})\\ \partial^{2}P_{\theta_{0}}(x_{1})&\partial^{2}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \partial^{n}P_{\theta_{0}}(x_{1})&\partial^{n}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{n}P_{\theta_{0}}(x_{i})\\ \end{pmatrix},\quad\quad\vec{x}=\begin{pmatrix}\tilde{\theta}_{1}\\ \tilde{\theta}_{2}\\ \vdots\\ \tilde{\theta}_{N}\end{pmatrix},\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix}.$ (71)
We point out that the matrix has dimensions $(n+1)\times N$ so if $n+1>N$ there are more equations than variables and a solution might not exist. Also, it is important to stress that the derivatives inside the matrix are all evaluated at point $\theta_{0}$ but we do not write it for the sake of clarity.
Rouché-Frobenius theorem states that the system has a solution if and only if $\operatorname{rank}(A|\vec{b})=\operatorname{rank}(A)$. If all the rows of $A$ are linearly independent matrix $(A|\vec{b})$ wouldn’t have a larger rank than $A$. Then, we are interested in the case where $A$ has linearly dependent rows. Notice that the first row is linearly independent of the others so the system $A\vec{x}=\vec{b}$ has no solution if and only if $A^{\prime}\vec{x}=\vec{\lambda}$ has no solution. From now on we study the latter system
$A^{\prime}=\begin{pmatrix}\partial P_{\theta_{0}}(x_{1})&\partial P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})\\ \partial^{2}P_{\theta_{0}}(x_{1})&\partial^{2}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \partial^{n}P_{\theta_{0}}(x_{1})&\partial^{n}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{n}P_{\theta_{0}}(x_{i})\\ \end{pmatrix},\quad\quad\vec{x}=\begin{pmatrix}\tilde{\theta}_{1}\\ \tilde{\theta}_{2}\\ \vdots\\ \tilde{\theta}_{N}\end{pmatrix},\quad\quad\vec{\lambda}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}.$ (72)
The matrix $A^{\prime}$ has dimensions $n\times N$ and since the last column is linearly dependent (is the sum of the $N-1$ first columns) there are at most $N-1$ linearly independent rows. As we mentioned before linearly dependent rows are a necessary condition for the system not having a solution. For that reason, we suppose that there is a row $m$ that can be written as,
$\begin{pmatrix}\partial^{m}P_{\theta_{0}}(x_{1})&\partial^{m}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{m}P_{\theta_{0}}(x_{i})\\ \end{pmatrix}=\sum_{l}\alpha_{l}\begin{pmatrix}\partial^{l}P_{\theta_{0}}(x_{1})&\partial^{l}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{l}P_{\theta_{0}}(x_{i})\\ \end{pmatrix}.$ (73)
Without losing generality we can consider that $m\neq 1$. Now we consider the $m$ row of the augmented matrix $(A^{\prime}|\vec{\lambda})$,
$\begin{pmatrix}\partial^{m}P_{\theta_{0}}(x_{1})&\partial^{m}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{m}P_{\theta_{0}}(x_{i})&0\\ \end{pmatrix}.$ (74)
If the $m$ row of the augmented matrix Eq. (74) is linearly independent from the rows of $(A^{\prime}|\vec{\lambda})$ then $\operatorname{rank}(A^{\prime}|\vec{\lambda})>\operatorname{rank}(A^{\prime})$ and there is no solution.
Thus we investigate whether it is linearly dependent or not. According to our assumption, we can write Eq. (74) as,
$\begin{pmatrix}\partial^{m}P_{\theta_{0}}(x_{1})&\dots&-\sum_{i}^{N-1}\partial^{m}P_{\theta_{0}}(x_{i})&0\\ \end{pmatrix}=\sum_{l}\alpha_{l}\begin{pmatrix}\partial^{l}P_{\theta_{0}}(x_{1})&\dots&-\sum_{i}^{N-1}\partial^{l}P_{\theta_{0}}(x_{i})&0\\ \end{pmatrix}$ (75)
but the first element of the sum,
$\alpha_{1}\begin{pmatrix}\partial P_{\theta_{0}}(x_{1})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})&0\\ \end{pmatrix}$ (76)
is not a row of the matrix $(A^{\prime}|\vec{\lambda})$ so Eq. (74) is linearly independent when $\alpha_{1}\neq 0$. Then, we conclude that the system given in Eq. (71) has no solution if and only if,
$\begin{pmatrix}\partial^{m}P_{\theta_{0}}(x_{1})&\partial^{m}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{m}P_{\theta_{0}}(x_{i})\\ \end{pmatrix}=\sum_{l}\alpha_{l}\begin{pmatrix}\partial^{l}P_{\theta_{0}}(x_{1})&\partial^{l}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{l}P_{\theta_{0}}(x_{i})\\ \end{pmatrix},$ (77)
where $\alpha_{1}\neq 0$.
Now we look at the Bhattacharyya bound
$(\Delta\hat{\Theta}_{\theta_{0}})^{2}\geq\max_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}\bm{C}\bm{a}}.$ (78)
We study the case when the BhB variance is unbounded. It is unbounded if and only if there is a vector $\bm{a}^{\prime}$ that satisfies $\bm{C}\cdot\bm{a}^{\prime}=0$ and $\bm{a}^{\prime\top}\bm{\lambda}\neq 0$. This vector $\bm{a}^{\prime}$ exists if and only if the next system of equations
$\begin{pmatrix}\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial P_{\theta_{0}}(x_{i})&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{2}P_{\theta_{0}}(x_{i})&\dots&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{2}P_{\theta_{0}}(x_{i})\partial P_{\theta_{0}}(x_{i})&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{2}P_{\theta_{0}}(x_{i})\partial^{2}P_{\theta_{0}}(x_{i})&\dots&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{2}P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{n}P_{\theta_{0}}(x_{i})\partial^{2}P_{\theta_{0}}(x_{i})&\dots&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{n}P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\\ \end{pmatrix}\begin{pmatrix}k\\ a_{2}\\ \vdots\\ a_{n}\end{pmatrix}=0,$ (79)
has a solution and also $k\neq 0$ because the condition that $\bm{a}^{\prime\top}\bm{\lambda}\neq 0$ has to be fulfilled too.
The system of equations can be rewritten as,
$\begin{pmatrix}\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{2}P_{\theta_{0}}(x_{i})&\dots&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{2}P_{\theta_{0}}(x_{i})\partial^{2}P_{\theta_{0}}(x_{i})&\dots&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{2}P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\\ \vdots&\ddots&\vdots\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{n}P_{\theta_{0}}(x_{i})\partial^{2}P_{\theta_{0}}(x_{i})&\dots&\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{n}P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\\ \end{pmatrix}\begin{pmatrix}a_{2}\\ \vdots\\ a_{n}\end{pmatrix}=k\begin{pmatrix}\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial P_{\theta_{0}}(x_{i})\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\end{pmatrix}.$
Rouché-Frobenius theorem states that the system has a solution if and only if $\operatorname{rank}(A|\vec{b})=\operatorname{rank}(A)$. This is equivalent to say that $\vec{b}$ is a linear combination of columns of the matrix $(A)$, i.e.
$\begin{pmatrix}\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial P_{\theta_{0}}(x_{i})\\ \vdots\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\end{pmatrix}=\sum_{l\neq 1}\alpha_{l}\begin{pmatrix}\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{l}P_{\theta_{0}}(x_{i})\partial P_{\theta_{0}}(x_{i})\\ \vdots\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{l}P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\end{pmatrix}.$
Rearranging the previous expression,
$\begin{pmatrix}\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{m}P_{\theta_{0}}(x_{i})\partial P_{\theta_{0}}(x_{i})\\ \vdots\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{m}P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\end{pmatrix}=\sum_{l=1,l\neq m}\alpha^{\prime}_{l}\begin{pmatrix}\sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{l}P_{\theta_{0}}(x_{i})\partial P_{\theta_{0}}(x_{i})\\ \vdots\\ \sum_{i}\frac{1}{P_{\theta_{0}}(x_{i})}\partial^{l}P_{\theta_{0}}(x_{i})\partial^{n}P_{\theta_{0}}(x_{i})\end{pmatrix}.$ (86)
where $\alpha_{1}^{\prime}\neq 0$. That is, there is a column $m$ that is a combination of the first column and other columns of $\bm{C}$ Note that the elements of the $l$ column of the rhs of Eq. (86) can be written as,
$\begin{pmatrix}\partial^{l}P_{\theta_{0}}(x_{1})&\partial^{l}P_{\theta_{0}}(x_{2})&\dots&\partial^{l}P_{\theta_{0}}(x_{N})\\ \end{pmatrix}.\begin{pmatrix}\frac{1}{P_{\theta_{0}}(x_{1})}\partial^{j}P_{\theta_{0}}(x_{1})\\ \frac{1}{P_{\theta_{0}}(x_{2})}\partial^{j}P_{\theta_{0}}(x_{2})\\ \vdots\\ \frac{1}{P_{\theta_{0}}(x_{N})}\partial^{j}P_{\theta_{0}}(x_{N})\\ \end{pmatrix},$
where $j$ goes from 1 to $n$ for each element of the column. Defining
$A_{m}$ $=\begin{pmatrix}\partial^{m}P_{\theta_{0}}(x_{1})&\partial^{m}P_{\theta_{0}}(x_{2})&\dots&\partial^{m}P_{\theta_{0}}(x_{N})\\ \end{pmatrix},$ (88)
$B_{j}$ $=\begin{pmatrix}\frac{1}{P_{\theta_{0}}(x_{1})}\partial^{j}P_{\theta_{0}}(x_{1})\\ \frac{1}{P_{\theta_{0}}(x_{2})}\partial^{j}P_{\theta_{0}}(x_{2})\\ \vdots\\ \frac{1}{P_{\theta_{0}}(x_{N})}\partial^{j}P_{\theta_{0}}(x_{N})\\ \end{pmatrix}.$ (93)
Equation (86) can be written as,
$A_{m}\cdot B_{j}$ $=\sum_{l=1}\alpha_{l}\>A_{l}\cdot B_{j}\quad\forall j.$ (94)
This implies
$A_{m}$ $=\sum_{l=1}\alpha_{l}\>A_{l}.$ (95)
Substituting the definitions given in Eq. (88) in Eq. (95) we get
$\begin{pmatrix}\partial^{m}P_{\theta_{0}}(x_{1})&\partial^{m}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{m}P_{\theta_{0}}(x_{i})\\ \end{pmatrix}=\sum_{l}\alpha_{l}\begin{pmatrix}\partial^{l}P_{\theta_{0}}(x_{1})&\partial^{l}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{l}P_{\theta_{0}}(x_{i})\\ \end{pmatrix},$ (96)
We have proved that BhB is unbounded if and only if Eq. (96) is true. We note that Eq. (96) is the same as Eq. (77). This is the same as saying that there is no estimator satisfying the conditions. As stated before, the matrix of Eq. (71) has at most $N$ linearly independent rows associated with $N$ independent conditions. Consider the case where the first $N-1$ higher order derivatives are linearly independent, then $n^{th}$order BhB bound will be equal to the $(N-1)^{th}$ BhB bound if $n\geq N-1$.
B.1 Bhattacharyya bounds for pure states
In this section, we show that contrary to what was recently claimed in the literature [30], Bhattacharyya bounds for pure states are computable and can be different from the Cramer-Rao bound in general. For this aim, we first explore why the conclusion exposed in [30], namely that for pure states Bhattacharyya bounds are the same as Cramer-Rao bound is faulted. After that, we present a simple counterexample of the conclusion. The proof in [30] is based on the fact that for pure states, $L_{i}$ (which represent the analogous of SLD but for higher derivatives) are bounded. From this, the authors concluded that the Bhattacharyya matrix is then filled with unbounded terms and when taking the optimization problem, one only looks at the first entry of the matrix which corresponds to the quantum Fisher Information. As exposed in [Yating], even when the operators $L_{i}$ are unbounded the product of $\rho L_{i}L_{j}$ may not be so, and this is what appears in the BhBattacarya matrix. The authors there argue that the usual Cramer-Rao bound (and for extension all Bhattacharyya bounds) is only constructed for bounded operators. For that reason, they “rederive” the Cramer-Rao bound for a newly defined operator $Q=L\sqrt{\rho}$. In the same fashion, one could do the same with Bhattacharyya bounds and write the Bhattacharyya matrix as a function of $Q_{i}$ operators. [RR: This we have to explicitly do here. ]This is shown in the next example. Consider a state $|\psi\rangle=\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$ where $x$ is the parameter one wants to estimate. From this state one can construct $Q_{i}$ operators that fulfill
$\frac{1}{2}(\sqrt{\rho}Q^{\dagger}_{j}+Q_{j}\sqrt{\rho})=\frac{\partial^{j}\rho}{\partial x^{j}},$ (97)
$\sqrt{\rho}Q_{j}-Q^{\dagger}_{j}\sqrt{\rho}=0,$ (98)
where the second equation comes from the hermiticity of $L_{j}$ [Yating]. After simple algebra one gets that (up to second order in Bhattacharyya bounds) in the basis $\{|\psi\rangle,|\psi^{\perp}\rangle\}$ the operators read
$Q_{1}=\begin{pmatrix}\langle\psi|\frac{\partial\rho}{\partial x}|\psi\rangle&0\\ 2\langle\psi^{\perp}|\frac{\partial\rho}{\partial x}|\psi\rangle&0\end{pmatrix},$ (99)
and
$Q_{2}=\begin{pmatrix}\langle\psi|\frac{\partial^{2}\rho}{\partial x^{2}}|\psi\rangle&0\\ 2\langle\psi^{\perp}|\frac{\partial^{2}\rho}{\partial x^{2}}|\psi\rangle&0\end{pmatrix},$ (100)
where $\rho=|\psi\rangle\langle\psi|$. From this, one can rewrite the quantum Bhattacharyya matrix (up to the second order) in terms of $Q_{j}$ as
$Q_{BhB}=\begin{pmatrix}\operatorname{Tr}{Q^{\dagger}_{1}Q_{1}}&\frac{1}{2}\left(\operatorname{Tr}{Q^{\dagger}_{1}Q_{2}}+\operatorname{Tr}{Q^{\dagger}_{2}Q_{1}}\right)\\ \frac{1}{2}\left(\operatorname{Tr}{Q^{\dagger}_{1}Q_{2}}+\operatorname{Tr}{Q^{\dagger}_{2}Q_{1}}\right)&\operatorname{Tr}{Q^{\dagger}_{2}Q_{2}}\end{pmatrix}.$ (101)
In this way, one can compute the second orther Bhattacharyya bound as the first element of the inverse of $Q_{BhB}$, i.e. $\Delta_{BhB}=Q^{-1}_{BhB,11}$. In Fig. (3) one can clearly see that the second order Bhattacharyya bound and CR bound differ.
Figure 3: Plot of the variances for different values of $x$. The upper one corresponds to the QBhB bound and the other to CR bound.
L_1 and L_2 formulas: Doing so, one gets
$L_{1}$ $=\left(\begin{array}[]{cc}4\theta(1-2\lambda)\sin\left(2\theta^{2}\right)&4i\theta(2\lambda-1)\cos\left(2\theta^{2}\right)\\ -4i\theta(2\lambda-1)\cos\left(2\theta^{2}\right)&4\theta(2\lambda-1)\sin\left(2\theta^{2}\right)\\ \end{array}\right),$ (104)
$L_{2}$ $=\left(\begin{array}[]{cc}\frac{4(2\lambda-1)\left(\theta^{2}\left(\cos\left(2\theta^{2}\right)-2\lambda+1\right)-(\lambda-1)\lambda\sin\left(2\theta^{2}\right)\right)}{(\lambda-1)\lambda}&\frac{4i(2\lambda-1)\left((\lambda-1)\lambda\cos\left(2\theta^{2}\right)+\theta^{2}\sin\left(2\theta^{2}\right)\right)}{(\lambda-1)\lambda}\\ -\frac{4i(2\lambda-1)\left((\lambda-1)\lambda\cos\left(2\theta^{2}\right)+\theta^{2}\sin\left(2\theta^{2}\right)\right)}{(\lambda-1)\lambda}&\frac{4(2\lambda-1)\left((\lambda-1)\lambda\sin\left(2\theta^{2}\right)-\theta^{2}\left(\cos\left(2\theta^{2}\right)+2\lambda-1\right)\right)}{(\lambda-1)\lambda}\\ \end{array}\right).$ (107)
From this, one can construct the BhBattacharyaa matrix which in this case reads
Appendix C Remarks to Ref. [30]
In this section, we raise some remarks to some claims made in Ref. [30]. Specifically, we present some arguments against the implication that if the matrix $\bm{Q}$ appearing in Eq. (14) is singular, the bound is divergent. Moreover, we also argue against the statement that for pure states all Bhattacharyya bounds give the same value as the CR bound. We make our arguments stronger using simple counterexamples of both propositions. Even though our results are only concerned with the Bhattacharyya bound, we think one could extend them for the other bounds presented in Ref. [30].
C.1 Invertibility of the matrix
In Ref. [30], Gessner and Smerzi argue that if the BhB matrix $\bm{C}$ is singular (non-invertible) the corresponding BhB bound will be divergent. While this may seem true at first inspection, we prove that is not true on a closer look. We now present our objections using the quantum version of the BhB bound, even though the conclusions also hold in the classical case. Moreover, we restrict ourselves to the second order. Let us first invoke the formula which gives the Bhattacharyya bound for the quantum case:
$(\Delta\hat{\Theta}_{\theta_{0}})^{2}\geq\max_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}\bm{Q}\bm{a}}=\bm{\lambda}^{\top}\bm{Q}^{-1}\bm{\lambda}.$ (108)
as explained in Ref. [30]. It is easy to see, that if the matrix is singular, one could always obtain zero in the denominator of Eq. (108). However, it does not follow that $\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}$ will be zero, since it may be the case that the eigenvector with null eigenvalue of $Q_{\mathrm{BhB}}$ may be orthogonal to $\bm{\lambda}$. This would imply that $\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}$ is zero over zero which is undefined so one cannot say that the variance is unbounded. In addition, we find illuminating to show the discrepancy with the aforementioned claim using a simple example. Consider we have a state given by $\rho=\theta|0\rangle\langle 0|+(1-\theta)|1\rangle\langle 1|$ where $\theta$ is the parameter we would like to estimate. It is clear that this state fulfills $\frac{\partial\rho}{\partial\theta}\neq 0$ and $\frac{\partial^{2}\rho}{\partial\theta^{2}}=0$. Then, $L_{2}=0$ and $L_{1}=1/\theta|0\rangle\langle 0|+1/(\theta-1)|1\rangle\langle 1|$ and the QBhB matrix will read
$\bm{Q}_{BhB^{2}}=\begin{pmatrix}\frac{1}{\theta(1-\theta)}&0\\ 0&0\\ \end{pmatrix},$ (111)
and
$\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}=\sup_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}Q_{\mathrm{BhB}}\bm{a}}=\sup_{\bm{a}}\frac{a^{2}_{1}}{\frac{a^{2}_{1}}{\theta(1-\theta)}}=\frac{1}{\frac{1}{\theta(1-\theta)}},$ (112)
which is essentially the QCRB, since the QFI is equal to $\frac{1}{\theta(1-\theta)}$. So, in this case, even if the QBhB matrix is singular, there exists an estimator that saturates the QBhB even though its value corresponds to the one given by the QCRB.
C.2 Bhattacharyya bounds for pure states
The literature claims that for pure states, Abel quantum bounds and in particular $n^{th}$order QBhB bounds are always equal to QCRB [30]. From our results, this would imply that there is always an estimator satisfying all the unbiasedness conditions. But in fact, it is not true. In the following, we give a simple counter-example and we point out why wrong conclusions are drawn in [30]. Consider the diagonal state $\rho=\lambda|0\rangle\langle 0|+(1-\lambda)|1\rangle\langle 1|$, where in the limit $\lambda\rightarrow 0$ it corresponds to a pure state. The state evolves under the unitary $U_{\theta}=e^{-i\theta\sigma_{x}}$. For this model the first and the third derivatives are equal, $\partial_{\theta}\rho_{\theta}=\partial^{3}_{\theta}\rho_{\theta}$ and
$\bm{Q}_{BhB^{3}}=\left(\begin{array}[]{ccc}4(1-2\lambda)^{2}&0&-16(1-2\lambda)^{2}\\ 0&-\frac{4(1-2\lambda)^{2}}{(\lambda-1)\lambda}&0\\ -16(1-2\lambda)^{2}&0&64(1-2\lambda)^{2}\\ \end{array}\right).$ (116)
Note that the first and the third rows are proportional. The bound in $\lambda\rightarrow 0$ (pure state), is given by
$\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}=\lim_{\lambda\rightarrow 0}\sup_{\bm{a}}\frac{\left(\bm{a}^{\top}\bm{\lambda}\right)^{2}}{\bm{a}^{\top}\bm{Q}_{\mathrm{BhB}}\bm{a}}.$ (117)
The supreme is achieved for $\bm{a}=(a_{1},0,\frac{a_{1}}{4})$ since it is an eigenvector of $\bm{Q}_{\mathrm{BhB}}$ with eigenvalue 00. In that case one has
$\left(\Delta\theta_{\mathrm{est}}\right)_{BhB}^{2}=\lim_{\lambda\rightarrow 0}\sup_{a_{1}}\frac{a_{1}}{0}=\lim_{\lambda\rightarrow 0}\infty=\infty$ (118)
In this example the BhB bound diverges while the quantum Fisher is finite –and hence the QCRB–, showing that for pure states the quantum Abel bounds and in particular the QBhB bounds can be different than the QCRB and one may be interested in computing them. The key is to consider the limit $\lambda\rightarrow 0$ after the supremum in Eq. (108). In the literature, the limit was taken before the supremum and it was observed that the matrix $\bm{Q}$ had some divergent elements that would not contribute to the bound.